OpenStudy (anonymous):

What is the equation of a line that passes through (7, -2) and has a slope of 3?

OpenStudy (anonymous):

Again, y = mx + b -2 = 3(7) + b b = -2 - 21 b = -23 Eqn: y = 3x - 23 :)